# Thompson Sampling for Repeated Newsvendor

### v1

@@ 3.2.1. Key Step 1: Regret Decomposition: Lipchitz Continuity
(5a) $\leq\max\{h,p\}\cdot\mathbb{E}\left[\sum_{t=1}^{T}\left|\mathbb{E}\left[y_{t}\right]-y_{\star}\right|\right]$
@@
Inequality (5a) comes from the following case discussion on $\min\{y_{t},D_{t}\}-\min\left\{y_{*},D_{t}\right\}$:
Case 1: $D_{t}>y_{t}$: In this case, $\min\left\{y_{t},D_{t}\right\}-\min\left\{y_{*},D_{t}\right\}=y_{t}-\min\left\{y_{*},D_{t}\right\}\geq y_{t}-y_{*}$. Then we have
$\mathbb{E}[h(y_{t}-y_{*})-(h+p)(\min\left\{y_{t},D_{t}\right\}-\min\left\{y_{*},D_{t}\right\})]$ $\leq-p\mathbb{E}[y_{t}-y^{*}]$
$=-p\mathbb{E}[\mathbb{E}[y_{t}]-y^{*}]$
$\leq p\mathbb{E}[\left|\mathbb{E}[y_{t}]-y^{*}\right|].$
Case 2: $D_{t}\leq y_{t}$: In this case $\min\left\{y_{t},D_{t}\right\}-\min\left\{y_{*},D_{t}\right\}=D_{t}-\min\left\{y_{*},D_{t}\right\}\geq 0$. Similarly,
$\mathbb{E}[h(y_{t}-y_{*})-(h+p)(\min\left\{y_{t},D_{t}\right\}-\min\left\{y_{*},D_{t}\right\})]$ $\leq h\mathbb{E}[y_{t}-y^{*}]$
$=h\mathbb{E}[\mathbb{E}[y_{t}]-y^{*}]$
$\leq h\mathbb{E}[\left|\mathbb{E}[y_{t}]-y^{*}\right|].$
@@
3.2.2. Key Step 2: Confidence Analysis of $\left|\mathbb{E}\left[y_{t}\right]-y_{\star}\right|$
Before we proceed, we give the definition for $y_{t}$ and $y_{\star}$ as follows:
Lemma 3.2.
The order quantity $y_{t}$ at $t$ and the optimal myopic order quantity $y_{t}$ satisfies
(6) $y_{*}(\theta_{\star})$ $=F^{-1}_{\theta_{\star}}(\frac{p}{p+h})=\frac{1}{\theta_{\star}}\left(-\ln(\frac{h}{p+h})\right)^{1/k}$
(7) $y_{t}(\theta_{t})$ $=F^{-1}_{\theta_{t}}(\frac{p}{p+h})=\frac{1}{\theta_{t}}\left(-\ln(\frac{h}{p+h})\right)^{1/k}$
where $F^{-1}$ is the inverse cumulative distribution function of the demand distribution. Moreover, $\mathbb{E}\left[\frac{1}{\theta_{t}}\right]=\frac{\beta_{t}}{\alpha_{t}-1}$.
By examining the expressions for $y_{t}$ and $y_{\star}$ in equations (7) and (6), we can directly derive that:
[…]
Lemma 3.3.
For each $t\in[T]$, with probability $\geq 1-\delta/T$, the realization of demand $D_{t}\sim\operatorname{Weibull}(\theta_{\star})$ will be in the range $[\underline{D},\overline{D}]$ such that,
$\underline{D}=\left(\frac{\ln{\left(\frac{2T}{2T-\delta}\right)}}{\theta_{\star}}\right)^{\frac{1}{k}},\qquad\overline{D}=\left(\frac{\ln{\left(\frac{2T}{\delta}\right)}}{\theta_{\star}}\right)^{\frac{1}{k}}.$
Lemma 3.3 is proved in Appendix A.1. This lemma ensures that, with high probability, the demand realizations are confined within the specified range, which is crucial for later analysis.
Next, we provide confidence bound for how close the $\frac{1}{\theta_{t}}$ and its mean $\frac{\beta_{t}}{\alpha_{t}-1}$ is. Ideally, as $t$ increases, $\theta_{t}$ will converge to $\theta_{*}$ and $\frac{1}{\theta_{t}}$ will converge to $\frac{\beta_{t}}{\alpha_{t}-1}$. The following lemma shows the rate of convergence as follows:
Lemma 3.4.
For any $t\in[T]$ and for any realization of $\theta_{\star}$,
$\mathbb{P}\left(\left|\frac{\beta_{t}}{\alpha_{t}-1}-\frac{1}{\theta_{\star}}\right|\geq\sqrt{\ln{\left(\frac{2t^{2}}{\delta}\right)}}\left(\overline{D}^{k}+\frac{2}{\theta_{\star}}\right)\sqrt{\frac{t}{\left(\alpha_{t}-1\right)^{2}}}\right)\leq\frac{\delta}{t^{2}}.$
Lemma 3.4 is proved in Appendix A.2. This lemma provides a probabilistic bound on the estimation error of $\frac{1}{\theta_{t}}$, which is key in assessing the accuracy of the order quantity decisions over time.
Combining these results, we can bound $\left|\mathbb{E}\left[y_{t}\right]-y_{\star}\right|$ as follows:
$\left|\mathbb{E}\left[y_{t}\right]-y_{\star}\right|$ $\leq\left(-\ln(\frac{h}{p+h})\right)^{1/k}\sqrt{\ln{\left(\frac{2t^{2}}{\delta}\right)}}\left(\overline{D}^{k}+\frac{2}{\theta_{\star}}\right)\sqrt{\frac{t}{\left(\alpha_{t}-1\right)^{2}}}$
$\leq\left(-\ln(\frac{h}{p+h})\right)^{1/k}\left(\overline{D}^{k}+\frac{2}{\theta_{\star}}\right)\sqrt{2\ln{\left(\frac{T}{\delta}\right)}}\sqrt{\frac{t}{\left(\alpha_{t}-1\right)^{2}}}.$
Lemma 3.5 ((Chuang and Kim, 2023)).
$\alpha_{t}$ $=\alpha_{0}+\sum_{i=0}^{t-1}\delta_{i}$
$=\alpha_{0}+\sum_{i=0}^{t-1}\mathbb{E}_{\theta_{\star}}^{\pi}\left[\delta_{i}\mid H_{i-1}\right]+\sum_{i=0}^{n-1}\left(\delta_{i}-\mathbb{E}_{\theta_{\star}}^{\pi}\left[\delta_{i}\mid H_{i-1}\right]\right)$
$=\alpha_{0}+\sum_{i=0}^{t-1}\left(1-e^{-\theta_{\star}y_{i}^{k}}\right)+M_{t}$
Where,
$M_{t}=\sum_{i=0}^{t-1}\left(\delta_{i}-\mathbb{E}\left[\delta_{i}\mid\mathcal{H}_{i-1}\right]\right)-1.$
$\mathbb{E}_{\theta_{\star}}^{\pi}$ denotes the expectation operator under admissible Bayesian policy $\pi\in\Pi$ given that the true unknown parameter is $\theta_{\star}\in\mathbb{R}_{+}$.
From the above lemma 3.5, we can see that as long as $y_{t}$ has a lower bound, we are able to derive the upper bound for regret.
3.2.3. Key Step 3: Uniform Lower Bound of $y_{t}$
In order to establish the regret bound, it is essential to establish a uniform lower bound for $y_{t}$, as this will play a crucial role in our subsequent derivations.
According to Lemma 3.6, we have:
Lemma 3.6.
The proof of Lemma 3.6 is provided in Appendix A.3.
Building upon this lemma, we proceed by conditioning on the event that $\frac{1}{\theta_{t}}>\frac{\beta_{t}}{2\alpha_{t}}$ and that the demand $D_{t}$ satisfies $D_{t}\geq\underline{D}$. Under these conditions, we can derive a lower bound for $y_{t}$ as follows:
(9a) $y_{t}$ $=\frac{1}{\theta_{t}}\left(-\ln(\frac{h}{p+h})\right)^{1/k}$
(9b) $\geq\frac{\beta_{t}}{2\alpha_{t}}\cdot\left(-\ln(\frac{h}{p+h})\right)^{1/k}$
(9c) $=\frac{1}{2}\left(-\ln(\frac{h}{p+h})\right)^{1/k}\cdot\frac{\beta_{0}+\sum_{i=1}^{t}\min\{y_{i},D_{i}\}^{k}}{\alpha_{0}+\sum_{i=1}^{t}\delta_{i}}$
(9d) $\geq\frac{1}{2}\left(-\ln(\frac{h}{p+h})\right)^{1/k}\cdot\min\left\{\frac{\beta_{0}}{\alpha_{0}},\underline{D}^{k}\right\}=L.$
Here, equation (9a) follows directly from the definition of $y_{t}$ as given in equation (7). Inequality (9b) utilizes the result from Lemma 3.6, indicating that with high probability, $\frac{1}{\theta_{t}}$ is bounded below by $\frac{\beta_{t}}{2\alpha_{t}}$. The equality in (9c) comes from the update rules for $\alpha_{t}$ and $\beta_{t}$ as defined in Algorithm 1. Finally, inequality (9d) is justified by applying Lemma 3.7, which is an auxiliary result crucial to our analysis.
Lemma 3.7.
From the closed-form expression (9c) and Lemma 3.7, we reveal the most important ascept of TS algorithm as follows:
  1. (1)
  2. (2)
    when $\delta_{i}=0$ (i.e. $D_{t}\geq y_{t}$). we get the censored demand, which indicates the past action is relatively small. Interestingly, since $\delta_{i}$ appears in the denominator in the closed-form expression (9c), TS naturally pushes future actions higher in subsequent periods, preventing the algorithm from getting stuck with poor estimates.
Applying Lemma 3.7 ( proved in Appendix A.4 ) in our context, and considering that $D_{t}\geq\underline{D}$, we conclude that $y_{t}$ is uniformly bounded below by $L$ for all $t$.
This uniform lower bound on $y_{t}$ is a critical to establish the regret bound. we plug back the lower bound to the definition of $\alpha_{t}$ in Lemma 3.5 to analyze the term $\left|\alpha_{t}-1\right|$, which analysis is referred in Appendix A.6.
To establish the regret, we use the technique of truncating $T$ and define a constant $C_{0}$ to encapsulate the terms independent of $t$ as follows:
Denote
$T_{0}=64\left(1-\exp\{-\theta_{\star}L^{k}\}\right)^{-2}\ln{\left(\frac{T}{\delta}\right)},$
define,
@@ 3.2.3. Key Step 3: Uniform Lower Bound of $y_{t}$
$C_{0}=\max\{h,p\}\left(-\ln(\frac{h}{p+h})\right)^{1/k}\left(\overline{D}^{k}+\frac{2}{\theta_{\star}}\right)\sqrt{2\ln{\left(\frac{T}{\delta}\right)}}.$
This allows us to express the regret bound more succinctly. Combining equation (5), section (3.2.2) and the discussions above, we can now bound the cumulative regret by truncating $T$ as follows:
$\operatorname{Regret(T,\theta_{\star})}$ $\leq C_{0}\sum_{t=1}^{T}\frac{\sqrt{t}}{\alpha_{t}-1}$
$\leq\begin{cases}C_{0}\cdot\left(T_{0}^{\frac{3}{2}}\cdot\frac{1}{\alpha_{0}-1}\right)&t\leq T_{0},\\ C_{0}\cdot\left(4\left(1-\exp\{-\theta_{\star}L^{k}\}\right)^{-1}\sqrt{T}\right)&t>T_{0}.\end{cases}$
3.2.4. Putting All together
In this section, we synthesize our previous findings to derive a comprehensive regret bound for the TS algorithm for the newsvendor problem.
$\operatorname{Regret(T,\theta_{\star})}$
[…]
@@ 3.2.4. Putting All together
$\leq C_{0}\cdot\left(512\left(1-\exp\{-\theta_{\star}L^{k}\}\right)^{-3}\cdot\ln\left(\frac{T}{\delta}\right)^{\frac{3}{2}}+4\left(1-\exp\{-\theta_{\star}L^{k}\}\right)^{-1}\sqrt{T}\right)\qquad\text{($y_{t}$ lower bound)}.$) .
As shown in the display, the three inequalities precisely correspond to the three key steps outlined in Section 3.2.1, Section 3.2.2, and Section 3.2.3.
3.3. Insights
In our analysis of the newsvendor problem with censored demand data, we employ Thompson Sampling (TS) to balance exploration and exploitation effectively. By modeling demand using a Weibull distribution with parameters estimated from prior data, TS updates these estimates as new sales data becomes available. In (9c), we come up with the uniform lower bound of action $y_{t}$, showing that large actions can enhance estimation accuracy, while small actions drive future TS-selected actions higher. In section 3.2.4, we conclude that our proof into three key steps. and we replace them by Assumptions 1 and 2 in Section 5, which enlightens us on how to extend the existing model to broader broader class of online learning.
4. Numerical Experiments
We conduct numerical experiments to evaluate the performance of TS in the repeated newsvendor problem and compare it against three benchmark policies. The first benchmark is the phased-UCB algorithm (Agrawal and Jia, 2019), which updates the confidence interval of the base-stock level at the beginning of each epoch, further subdividing each epoch into consecutive time steps. We denote this policy as UCB. The second benchmark is the non-parametric adaptive policy proposed by (Huh and Rusmevichientong, 2009), which generates ordering decisions dynamically over time. We denote this policy as OCO. Finally, we compare TS with the myopic policy from (Besbes et al., 2022). which is a deterministic policy where the decision-maker optimizes inventory decisions one period at a time, solving a single-period problem without considering how the chosen order quantity impacts future learning of demand parameters.
To systematically analyze the impact of different service levels, we define the service level as $\gamma=\frac{p}{p+h}$, where we fix $p=1$ and vary $h$ to achieve service levels of $50\%$, $90\%$, and $98\%$. For each experiment, we simulate the TS algorithm and the three benchmark policies on a common problem instance. Each algorithm is run for $100$ independent trials to mitigate randomness, and we report the average cumulative regret. We set the prior parameters of the Weibull distribution to $\alpha_{0}=\beta_{0}=4$ and consider a time horizon of $T=600$.
We present our results in two sets of plots:
  1. (1)
    Comparison of TS, UCB, and OCO. We plot the average cumulative regret of TS, UCB, and OCO to assess their relative learning performance in Figure 1.
  2. (2)
    Comparison of TS and Myopic. We compare the average cumulative regret of TS and the myopic policy against the optimal cost in Figure 2.
Our results demonstrate that TS consistently outperforms UCB and OCO across all service levels. Additionally, when comparing TS to the myopic policy, we observe that TS converge faster than Myopic, further reinforcing its effectiveness in balancing exploration and exploitation in the newsvendor setting.
Figure 1. Compare TS with OCO and UCB
Figure 2. Compare TS with Myopic Policy
5. Extensions to Online Learning with Censored Feedback

### v2

@@ 3.2.1 Key Step 3: Regret Decomposition via Lipschitz Continuity
$\leq\max\{h,p\}\cdot\mathbb{E}\left[\sum_{t=1}^{T}\left|\mathbb{E}\left[y_{t}\right]-y_{\star}\right|\right]$ (6a)
@@
Inequality (6a) comes from the following case discussion on $\min\{y_{t},D_{t}\}-\min\left\{y_{\star},D_{t}\right\}$:
Case 1: $D_{t}>y_{t}$: In this case, $\min\left\{y_{t},D_{t}\right\}-\min\left\{y_{\star},D_{t}\right\}=y_{t}-\min\left\{y_{\star},D_{t}\right\}\geq y_{t}-y_{\star}$. Then $\mathbb{E}[h(y_{t}-y_{\star})-(h+p)(\min\left\{y_{t},D_{t}\right\}-\min\left\{y_{\star},D_{t}\right\})]\leq-p\mathbb{E}[y_{t}-y_{\star}]=-p\mathbb{E}[\mathbb{E}[y_{t}]-y_{\star}]\leq p\mathbb{E}[\left|\mathbb{E}[y_{t}]-y_{\star}\right|]$.
Case 2: $D_{t}\leq y_{t}$: In this case $\min\left\{y_{t},D_{t}\right\}-\min\left\{y_{\star},D_{t}\right\}=D_{t}-\min\left\{y_{\star},D_{t}\right\}\geq 0$. Similarly, $\mathbb{E}[h(y_{t}-y_{\star})-(h+p)(\min\left\{y_{t},D_{t}\right\}-\min\left\{y_{\star},D_{t}\right\})]\leq h\mathbb{E}[y_{t}-y_{\star}]=h\mathbb{E}[\mathbb{E}[y_{t}]-y_{\star}]\leq h\mathbb{E}[\left|\mathbb{E}[y_{t}]-y_{\star}\right|]$.
@@
3.3 Putting All Together
@@
Denote $C_{0}=\max\{h,p\}\left(-\ln(\frac{h}{p+h})\right)^{1/k}\left(\overline{D}^{k}+\frac{2}{\theta_{\star}}\right)\frac{1}{k}\cdot\min\left\{L,\frac{1}{\theta_{*}}\right\}^{1/k-1}\sqrt{2\ln{\left(\frac{T}{\delta}\right)}}$.
Altogether, we have
$\operatorname{Regret(T,\theta_{\star})}$ $\leq\max\{h,p\}\cdot\sum_{t=1}^{T}\mathbb{E}\left[\left|\mathbb{E}\left[y_{t}\right]-y_{\star}\right|\right]\qquad\text{(Lipchitz Continuity)}$
@@
$\leq C_{0}\cdot\left(512\left(1-\exp\{-\theta_{\star}L^{k}\}\right)^{-3}\cdot\ln\left(\frac{T}{\delta}\right)^{\frac{3}{2}}+4\left(1-\exp\{-\theta_{\star}L^{k}\}\right)^{-1}\sqrt{T}\right)\qquad\text{($y_{t}$ lower bound)}.$
As shown, the three inequalities correspond precisely to the three key steps in Section 3.1, Section 3.2, and Section 3.2.1.
4 Extension beyond Weibull Distribution
In this section, we provide analysis for Bayesian regret for arbitrary demand distribution. We first denote $\hat{F}_{t}$ to be estimated CDF of demand and $F_{\star}$ be the true demand CDF for unknown arbitrary demand distribution.
Assumption 1.
There exists a parameter $\theta_{\star}\in\Theta$ such that $F_{\theta_{\star}}=F_{\star}$.
Assumption 2.
For the estimator $\hat{\theta}$ of the unknown parameter $\theta_{\star}$, we assume $|\hat{\theta}-\theta_{\star}|\leq C_{1}\cdot\|F_{\hat{\theta}}-F_{\star}\|_{\infty}$,
Assumption 3.
The newsvendor cost function satisfies $|g_{\theta_{1}}(y,D)-g_{\theta_{2}}(y,D)|\leq C_{2}\cdot|\theta_{1}-\theta_{2}|$.
Combing Assumption 2 with Assumption 3, we can get $|g_{\theta_{1}}(y,D)-g_{\theta_{2}}(y,D)|\leq C_{1}C_{2}\cdot\|F_{\theta_{1}}-F_{\theta_{2}}\|_{\infty}$.
4.1 Kaplan-Meier Estimator and Confidence Interval
Definition 3.
Huh et al. (2011)
$1-\hat{F}_{t}(x)=\prod_{s:Y_{(s)}\leqslant x}\left(\frac{t-s}{t-s+1}\right)^{\delta_{(s)}}$
Notice that for $1\leq s\leq t\leq T$, $Y_{(s)}$ and $\delta_{(s)}$ denote order statistics. Specifically, consider $T$ observations $\left\{\left(Y_{1},\delta_{1}\right),\ldots,\left(Y_{T},\delta_{T}\right)\right\}$, some of which are possibly censored (i.e., with $\delta_{t}=0$). To construct the KM estimator, we order the $T$ observations from smallest to largest according to the value of $Y_{t}$ ’s. Ties are broken by placing uncensored observations ($\delta_{t}=1$) before censored ones ($\delta_{t}=0$).
Definition 4 (Greedy Plug-in Estimator).
At each time $t$, we define a greedy parameter estimate $\hat{\theta}_{t}$ that minimizes the distance to the current empirical distribution $\hat{F}_{t}$ under KM estimator , and the corresponding plug-in estimator $F_{\hat{\theta}_{t}}$ such that $\|F_{\hat{\theta}_{t}}-\hat{F}_{t}\|_{\infty}\leq\|F_{\theta_{\star}}-\hat{F}_{t}\|_{\infty}$.
Lemma 7.
Bitouzé et al. (1999) According to Theorem 1 of Bitouzé et al. (1999) we have, for each $t\in[T]$ and let $\hat{F}_{t}$ be the Kaplan-Meier estimator of the distribution function $F$. And $G$ is the censoring distribution function. There exists an absolute constant $C$ such that, for any positive $\lambda$,
$\mathbb{P}\left(\sqrt{t}\left\|(1-G)\left(\hat{F}_{t}-F\right)\right\|_{\infty}>\lambda\right)\leqslant 2.5\mathrm{e}^{-2\lambda^{2}+C\lambda}.$
Lemma 8.
For each $t\in[T]$ and arbitrary x and $0<G(x)<1$. and set $\epsilon=\sqrt{\frac{1}{2t}\ln\left(\frac{1}{\delta}\right)}$, we have
$\|\hat{F}_{t}(x)-F(x)\|_{\infty}\leq\frac{\epsilon}{1-G(x)}.$
4.2 Regret Analysis
Theorem 2 (Bayesian Regret for general demand).
Under Assumptions 1 and 2, and with the KM-based plug-in estimator defined in Definition 4, the Bayesian regret of the policy $\pi$ over $T$ periods satisfies the following bound with probability at least $1-\delta$:
$\operatorname{BayeRegret}(T,\pi)$ $\leq\frac{16L(h+p)}{1-G_{\max}}\sqrt{T\ln(1/\delta)},$
where $G_{\max}<1$ is an upper bound on the CDF $G(x)$ over the demand support.
The complete proof for Theorem 2 is in Appendix A.8. We provide the three key steps here.
Step 1: Establishing the Plug-in Estimator Properties
By definition 4, Lemma 8 and assumption 1, we have $\|F_{\hat{\theta}_{t}}-\hat{F}_{t}\|_{\infty}\leq\|F_{\theta_{\star}}-\hat{F}_{t}\|_{\infty}=\|F_{\star}-\hat{F}_{t}\|_{\infty}$. This gives us $\|F_{\hat{\theta}_{t}}-F_{\star}\|_{\infty}\leq 2\|\hat{F}_{t}-F_{\star}\|_{\infty}$. Therefore $|\hat{\theta}_{t}-\theta_{\star}|\leq L\|F_{\hat{\theta}_{t}}-F_{\star}\|_{\infty}$. We define cost functions $g_{\hat{\theta}_{t}}(y_{t},D_{t})$, $\hat{g}(y_{t},D_{t})$, and $g_{\theta_{\star}}(y_{t},D_{t})$ for the estimator, KM estimator $\hat{\theta_{t}}$, and true parameter respectively. This yields $\left|g_{\hat{\theta}_{t}}(y_{t},D_{t})-\hat{g}(y_{t},D_{t})\right|\leq\left|g_{\theta_{\star}}(y_{\star},D_{t})-\hat{g}(y_{t},D_{t})\right|=\left|g(y_{\star},D_{t})-\hat{g}(y_{t},D_{t})\right|.$ According to Pilz (1991), the Bayesian estimator is the minimum mean square error estimator of posterior mean. We use KM estimator to upper bound the Bayesian regret per-period.
Step 2: UCB Regret Decomposition
We decompose the per-period Bayesian regret $g(y_{t},D_{t})-g(y_{\star},D_{t})$ as $g(y_{t},D_{t})-U_{t}(y_{\star})]+[U_{t}(y_{t}^{km})-g(y_{\star},D_{t})]$ , here we define an upper confidence bound sequence in Russo and Van Roy (2014) $U_{t}(y)=\hat{g}(y,D_{t})$ is the empirical cost based on the KM estimator and $y_{t}^{km}=\arg\min_{y}U_{t}(y)$ is the KM-optimal decision.
Step 3: Putting All Together
Lemma 9.
(Lipschitz Property) The newsvendor cost function satisfies $|g_{\theta_{1}}(y,D)-g_{\theta_{2}}(y,D)|\leq(h+p)\cdot\|F_{\theta_{1}}-F_{\theta_{2}}\|_{\infty}$.
We exploit the Lipschitz lemma to convert the KM estimation error $|\hat{F}_{t}(x)-F_{\star}(x)|\leq\frac{\epsilon_{t}}{1-G(x)}$ directly into bounds on cost function differences. For Term I, this gives $g(y_{t},D_{t})-U_{t}(y_{\star})\leq\frac{L_{1}\epsilon_{t}}{1-G(y_{\star})}$, and for Term II, we get $U_{t}(y_{t}^{km})-g(y_{\star},D_{t})\leq\frac{L_{2}\epsilon_{t}}{1-G_{\max}}$. Assuming that ($G(x)\leq G_{\max}<1$. Combining both terms and summing over time with $\epsilon_{t}=\sqrt{\frac{\ln(1/\delta)}{2t}}$ yields the final regret bound of $\frac{16L(h+p)}{1-G_{\max}}\sqrt{T\ln(T)}$.
5 Numerical Experiments
We conduct numerical experiments to evaluate the performance of Thompson Sampling (TS) in the repeated newsvendor problem, comparing it against three benchmarks. The first is the phased-UCB algorithm (Agrawal and Jia, 2019), which updates the confidence interval at the start of each epoch (UCB). The second is the non-parametric adaptive policy from Huh and Rusmevichientong (2009), which dynamically adjusts orders over time (OCO). The third is the deterministic myopic policy from Besbes et al. (2022), which optimizes single-period inventory decisions without accounting for future learning. To examine the impact of different service levels, defined as $\gamma=\frac{p}{p+h}$, we fix $p=1$ and vary $h$ to achieve service levels of 50%, 90%, and 98%. All policies are tested on a common problem instance with prior parameters $\alpha_{0}=\beta_{0}=4$ and horizon $T=600$. Each algorithm is run over 100 independent trials, and we report average cumulative regret.Results are shown in two plots: (1) a comparison of TS, UCB, and OCO (Figure 1), and (2) a comparison of TS and the myopic policy against the optimal cost (Figure 2). Across all service levels, TS consistently outperforms the benchmarks and converges faster than the myopic policy, highlighting its effectiveness in balancing exploration and exploitation.
Figure 1: Compare TS with OCO and UCB
Figure 2: Compare TS with Myopic Policy
6 Conclusions
We present the first systematic study on applying Thompson Sampling (TS) to the repeated newsvendor problem and provide an initial exploration of how our analytical framework can be extended to broader online learning problems with censored feedback. We establish frequentist regret bounds and offer insights into how TS automatically balances the trade-off between “large exploration” and “optimal exploitation.” Our analysis follows three key steps, which naturally generalize to broader settings. This work opens up a range of compelling research directions.
We extend the newsvendor problem to general online learning with censored feedback in Appendix A.1. We illustrate how our analytical framework naturally extends to broader settings of online learning with censored feedback, making it applicable to a wide range of problems where the feedback is censored.This extension requires only that the regret is Lipschitz continuous with respect to actions and that the relationship between the optimal action and the underlying parameter is continuous and monotone. Lastly, we conduct additional experiments in Appendix A.10.
Appendix A Appendix
A.1 Extensions to Online Learning with Censored Feedback